Determine the pH during the titration of 18.3 mL of 0.311 M hydrobromic acid by 0.153 M barium hydroxide at the following points: (1) Before the addition of any barium hydroxide (2) After the addition of 9.30 mL of barium hydroxide (3) At the equivalence point (4) After adding 22.3 mL of barium hydroxide
Answer #1

Answer:-

This question is answered by using simple concept of titration of strong acid and strong base.

The answer is given in the image,

2 Answer :- CHBYJ= 0.311M CBalon),] 20-153M ☺ (H307) = [HBr] = 0.311M - PH2-log(43 ot] = -log(0-311) Pu= 0. 51 9.30m2 BacoH):= 6. 8238х10 3 mv1. Excel Flobes of om = 6-82 38х16 3- с. 69јзхь 3 - 1-1325xip mo). (би)= 1-132 5 x 10 3х) обо (18•3+ 22.3)